Question 6 (1 point) Fill in the blanks. Pinterest claims that 0.27 of their app users are men. In a sample of 72 randomly chosen app users, around Assume each pick is independent. of them will be men, give or take 0 1) 19.44,14.20 2) 72,3.767 3) 19.44,0.27 )3.767,19.44 5) 19.44,3.767 4
Answer #1

n = 72 , p = 0.27

mean = np = 72*0.27 = 19.44

sd = sqrt(npq) = sqrt(72*0.27 * 0.73)= 3.767

option 5) is correct